4.Suppose that α=2i−3j+k and \eta=7i−5j+k , find a unit vector perpendicular to α
and \(\eta) respectively.
\\(\\frac{-2i+5j+11k }{5\\sqrt(6)}\\)
\\(\\frac{2i-5j+11k }{5\\sqrt(6)}\\)
\\(\\frac{2i+5j-11k }{5\\sqrt(6)}\\)
\\(\\frac{2i+5j+11k }{5\\sqrt(6)}\\)
1
Expert's answer
2020-05-20T19:51:49-0400

"\alpha =2i-3j+k, \ \ \beta =7i-5j+k"

The cross product "\alpha \times \beta" is a vector that is perpendicular to both "\alpha" and  "\beta" .

"\gamma =\alpha \times \beta =\begin{vmatrix}
i&j&k
\\ 2&-3&1\\7&-5&1
\end{vmatrix}=(-3\times 1-(-5)\times 1)i-(2\times 1-7\times 1)j+(2\times (-5)-7(-3))k=2i+5j+11k"


"|\gamma|=\sqrt{2^2+5^2+11^2}=\sqrt{150}=5\sqrt6"

Unit vector is "\frac{\gamma}{|\gamma|}" .


Answer: "\frac{2i+5j+11k}{5\sqrt6}"  a unit vector perpendicular to "\alpha" and "\beta" respectively.